OpenStudy (anonymous):

Find the value of x in each of the following equations. a)2^x-1=13 @ganeshie8

ganeshie8 (ganeshie8):

time to use latex @MARC_ :) is it \(2^{x-1} = 13\) or \(2^x-1 = 13\) ?

OpenStudy (anonymous):

\[2^{x-1}=13\] alright :)

ganeshie8 (ganeshie8):

looks good, here \(x-1\) is called the `logarithm` of \(13\) in base \(2\)

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

\[2^{x-1} = 13\] is equivalent to \[\large x-1 = \log_2 13\]

ganeshie8 (ganeshie8):

add 1 both sides and you're done : \[\large x = 1 + \log_2 13 \]

ganeshie8 (ganeshie8):

wolfram says \[x \approx 4.7\]

ganeshie8 (ganeshie8):

Alternatively you may simply use the log properties to solve \(x\) : \[2^{x-1} = 13\] take \(\ln\) both sides : \[ \ln (2^{x-1}) = \ln (13)\] by using the log property \(\color{Red}{\ln(a^x) = x*\ln(a)}\) : \[ (x-1)*\ln (2) = \ln (13)\] \[ (x-1) = \dfrac{\ln (13)}{\ln(2)}\] \[x = 1+\dfrac{\ln (13)}{\ln(2)}\]

OpenStudy (anonymous):

okay i get :)

OpenStudy (anonymous):

Thnx @ganeshie8

ganeshie8 (ganeshie8):

good :) remember that `logarithm` is same as the `exponent` \[\large a^{\color{Red}{x}} = y\] here \(\color{Red}{x}\) is called the \(\color{Red}{\text{logarithm}}\) of \(y\) in base \(a\)

ganeshie8 (ganeshie8):

example, we know : \[\large 2^{\color{Red}{3}} = 8\] here \(\color{Red}{3}\) is called the \(\color{Red}{\text{logarithm}}\) of \(8\) in base \(2\)

ganeshie8 (ganeshie8):

\[ \log_2 8=\color{Red}{3} \]

ganeshie8 (ganeshie8):

logarithm = exponent

OpenStudy (anonymous):

Thnx for ur explaination :)

ganeshie8 (ganeshie8):

yw!